OpenStudy (anonymous):

If f(x)= -4x^2+12x-9 and g(x)= -9-3x+x^2,find 6f(1/2)-2g(2). Please show me step by step.

jimthompson5910 (jim_thompson5910):

f(x)= -4x^2+12x-9 f(1/2)= -4(1/2)^2+12(1/2)-9 f(1/2)= -4(1/4)+12(1/2)-9 f(1/2)= -4/4+12/2-9 f(1/2)= -1+6-9 f(1/2)= -4 ------------------------------------------------------- g(x) = -9-3x+x^2 g(2) = -9-3(2)+(2)^2 g(2) = -9-3(2)+4 g(2) = -9-6+4 g(2) = -11 ------------------------------------------------------- 6*f(1/2) - 2*g(2) 6*(-4) - 2*(-11) -24 + 22 -2 So 6*f(1/2) - 2*g(2) = -2

OpenStudy (anonymous):

First off: You have to plug 1/2 into f(x). f(1/2)= -4*(1/2)^2+12(1/2)-9 Then, you need to plug 2 into g(x): g(2)=-9-3(2)+(2)^2 Next you need to put everything together. 6*[f(1/2)= -4*(1/2)^2+12(1/2)-9]-2*[g(2)=-9-3(2)+(2)^2] So you solve what's inside the parentheses And then you multiply what ever you get inside by the numbers outside of the parentheses And finally you subtract the numbers. And your final answer is -2